### v1

@@ Abstract
The accurate segmentation of coronary Digital Subtraction Angiography (DSA) images is essential for diagnosing and treating coronary artery diseases. Despite advances in deep learning-based segmentation, challenges such as low contrast, noise, overlapping structures, high intra-class variance, and class imbalance limit precise vessel delineation. To overcome these limitations, we propose the MSA-UNet3+: a Multi-Scale Attention enhanced UNet3+ architecture for coronary DSA image segmentation. The framework combined Multi-Scale Dilated Bottleneck (MSD-Bottleneck) with Contextual Attention Fusion Module (CAFM), which not only enhances multi-scale feature extraction but also preserve fine-grained details, and improve contextual understanding. Furthermore, we propose a new Supervised Prototypical Contrastive Loss (SPCL), which combines supervised and prototypical contrastive learning to minimize class imbalance and high intra-class variance by focusing on hard-to-classified background samples. Experiments carried out on a private coronary DSA dataset demonstrate that MSA-UNet3+ outperforms state-of-the-art methods, achieving a Dice coefficient of 87.73%, an F1-score of 87.78%, and significantly reduced Average Surface Distance (ASD) and Average Contour Distance (ACD). The developed framework provides clinicians with precise vessel segmentation, enabling accurate identification of coronary stenosis and supporting informed diagnostic and therapeutic decisions. The code will be released at the following GitHub profile link https://github.com/rayanmerghani/MSA-UNet3plus.
@@ 1 Introduction
Figure 1: Illustration of the desired characteristics of an encoder, SCE optimizes the embedding space by minimizing the distance between similar foreground samples (in blue) and maximizing the distance between dissimilar ones. PCL focuses on learning prototypes for foreground samples (in blue star), pulling these samples close to their respective prototypes while pushing hard negative instances (those close to the prototypes) further away.
@@
Segmentation networks in the reviewed studies, particularly encoder-decoder architectures, face critical limitations. Encoders often prioritize performance over semantic embedding [14], failing to cluster similar class vectors closely in the embedding space. This challenge intensifies in coronary DSA segmentation due to the background class’s structural diversity and ambiguous boundaries, causing significant intra-class variance and class imbalance. Current approaches learn discriminative representations for both classes but struggle with the negative class’s extreme diversity. Most methods also neglect hard negative samples—the most optimization-informative yet least abundant because of visual dissimilarity.
To address these limitations, we propose a deep learning framework combining prototype learning [15] and supervised contrastive loss [16]. Our hybrid approach (Fig. 1) improves encoder discriminability by: (1) clustering positive features while distancing them from negatives via supervised contrastive loss, and (2) refining background embeddings through prototypical contrastive loss that isolates hard negatives by prototype distance. This dual strategy explicitly targets class imbalance and intra-class variance by focusing on challenging samples.
@@
Figure 2: Architecture of the proposed MSA-UNet3+ model. The model integrates a Multi-Scale Dilated Bottleneck (MSD-Bottleneck) for multi-scale feature extraction and a Contextual Attention Fusion Module (CAFM) for enhanced contextual understanding. The M-encoder employs convolutional and transposed convolutional layers, while the decoders reconstruct the segmentation mask. This architecture enables precise segmentation of coronary arteries in DSA images by capturing both fine-grained details and broader structural information.
@@
We integrate the proposed Supervised Prototypical Contrastive Loss (SPCL) within our Multi-Scale Attention-modified UNet3+ (MSA-UNet3+) framework (Fig. 2). The architecture enhances UNet3+ through three key modifications: (1) a Multi-Scale Attention Encoder (M-encoder) replacing the traditional encoder for hierarchical feature extraction, (2) a Multi-Scale Dilated Bottleneck (MSD-Bottleneck) with Atrous Spatial Pyramid Pooling (ASPP) for multi-contextual feature fusion, and (3) a Contextual Attention Fusion Module (CAFM) that performs channel-wise feature recalibration. The decoder then synthesizes these refined multi-scale features to generate precise coronary DSA segmentation masks. The key contributions of this work can be summarized as follows.
  1. 1.
    We introduce a new hybrid loss function, Supervised Prototypical Contrastive Loss (SPCL), which is a hybrid of supervised contrastive loss and prototypical contrastive loss. The SPCL can mitigate the problem of class imbalance and high intra-class variance, which significantly enhances the encoder’s ability to discriminate the complex classes in the segmentation of coronary DSA images. This novel solution mitigates major drawbacks of current techniques, yielding improved segmentation performance.
  2. 2.
    Our method uniquely emphasizes and handles hard-to-classify background samples, which are generally ignored by common methods. Focusing on such difficult cases, the proposed framework maintains higher learning efficiency, more robust segmentation accuracy, and more applicable knowledge of how to treat class imbalance on coronary DSA images. This approach ensures improved generalization and performance compared to the implementation of actual clinical dynamics.
  3. 3.
    We develop and propose the Multi-Scale Attention-modified U-Net3+ (MSA-UNet3+), which is a general segmentation framework with a multi-scale attention encoder, a Multi-Scale Dilated Bottleneck and Contextual Attention Fusion Modules (CAFM) for reaching state-of-the-art performance. This architecture is specifically optimized for coronary DSA images and therefore helps improve feature extraction, contextual understanding, and accurate delineation of complex anatomical structures. We call our new modelnMSA-UNet3+, which advances the previous state-of-the-art in segmentation accuracy and computational efficiency in medical images.
@@ 2 Related Work
Medical image segmentation is critical for computer-aided diagnosis and treatment planning [17, 18]. While U-Net and its variants [19, 20] excel at capturing spatial-contextual information. However, standard U-Net struggles with fine details and class imbalance. Variants like Res-UNet and Attention Res-UNet address these issues. Recent advances include: (1) Convolutional Neural Networks (CNNs) based architectures for robust feature extraction [18, 21, 22], (2) Fully Convolutional Networks (FCNs) for flexible input sizes [5], (3) Recurrent Neural Networks (RNNs) /ConvLSTMs for temporal modeling [23, 24], and (4) Generative Adversarial Networks (GANs) for data augmentation [25]. Transformers (e.g., TransUNet [26]) and foundation models like SAM [20] now enable attention-based long-range dependencies. However, class imbalance remains a fundamental challenge, particularly in medical imaging where background dominance biases model performance [27].
@@ 2.1 Traditional Approaches to Class Imbalance
Traditional methods for handling class imbalance in machine learning, such as data resampling techniques, can also be applied to medical image segmentation. Oversampling techniques, such as creating synthetic samples from the minority class, aim to balance the class distribution [28]. However, naive oversampling methods can lead to overfitting, especially in the context of medical images, where generating truly representative synthetic data is challenging. Undersampling, on the other hand, involves removing samples from the majority class [29]. Although this can reduce computational cost and mitigate overfitting, it leads to information loss and potentially discarding valuable data. The effectiveness of these techniques is highly dependent on the specific dataset and the nature of the class imbalance [29]. However, even with oversampling, careful consideration must be given to the methods used to generate synthetic data to ensure that they accurately represent the underlying data distribution and prevent overfitting [28].
@@ 2.2 Loss Function Modifications
Addressing class imbalance directly within the loss function is another common strategy. While standard cross-entropy, is sensitive to class imbalance thus favors majority classes [30, 31], three key adaptations exist: (1) Weighted cross-entropy re-balances class contributions through manual weighting, preventing the model from being dominated by the majority class [31, 32], However, determining appropriate weights can be challenging and requires potentially hyperparameter tuning [33]; (2) Focal loss [30, 34] reduces the contribution of easily classified samples (often from the majority class) and focuses learning on more challenging samples from the minority classes; and (3) Tversky loss [35], a generalization of the Dice loss, offers another avenue to address class imbalance by adjusting the weights assigned to false positives and false negatives, allowing for a more balanced trade-off between precision and recall. Selection between losses depends on the dataset characteristics and desired precision-recall tradeoffs [30].
@@ 2.3 Advanced Techniques
Advanced techniques beyond traditional and loss function-based approaches have been explored to address class imbalance in medical image segmentation. Semi-supervised learning, using both labeled and unlabeled data, shows promise [36, 37]. Methods like contrastive learning and consistency regularization [38, 39] learn robust representations from limited labeled data, enhancing performance in minority classes. These approaches often involve data augmentation and enforcing consistency between predictions on augmented versions of the same image. For example, a multitask contrastive learning framework has been proposed to address class imbalance through global and local contrastive learning and multiscale uncertainty estimation [38]. Additionally, a dual-path framework tackles intra-class imbalance by separating objects into subclasses and using weighted maps to guide segmentation [40].
Although most contrastive learning studies focus on unsupervised settings, Khosla et al. [41] extended it to supervised learning by treating all samples from the same class as positive. Contrastive learning has also been applied to semantic segmentation. Chaitanya et al. [42] introduced techniques to exploit structural similarities in volumetric medical images within a self-supervised framework, categorizing slices of similar volumes as positive pairs and dissimilar ones as negative pairs. Similarly, Zhao et al. [43] proposed a fine-tuning approach using contrastive loss for semantic segmentation. However, these studies primarily aim to improve performance with limited data, rather than training segmentation encoders to inherently learn semantic embeddings.
Despite significant progress, there are challenges to address the class imbalance in medical image segmentation. Generating high-quality synthetic data remains difficult, and the optimal choice of loss functions or augmentation strategies is often dataset-specific [44]. Overcoming these challenges is critical for accurate coronary DSA image segmentation. This study integrates supervised and prototypical contrastive learning strategies, proposing a novel hybrid loss that combines supervised contrastive loss with prototypical contrastive loss. This approach enhances the encoder’s ability to generate discriminative features, enabling better differentiation between diverse classes. By focusing on challenging negative samples, the hybrid loss effectively addresses class imbalance and high intra-class variance.
@@ 3.1 Overview
In this study, we propose a Multi-Scale Attention modified UNet3+ (MSA-UNet3+) network that incorporates contrastive learning for coronary DSA image segmentation in CAD diagnosis. Existing methods face significant challenges in processing coronary DSA images due to two interrelated factors: the ambiguous background class containing diverse anatomical structures (ribs, spine, diaphragm, and lungs) and the consequent high intra-class variance and class imbalance. To address these limitations, we introduce the Supervised Prototypical Contrastive Loss (SPCL), which unifies the advantages of supervised and prototypical contrastive learning. The supervised component improves discriminative feature learning, while the prototypical component optimizes class-aware embeddings. Through this dual mechanism, SPCL significantly improves foreground-background separation while effectively addressing class imbalance through prototype-based hard sample mining.
@@ 3.2 Supervised Prototypical Contrastive Learning
Fig. 1 illustrates the pipeline of our proposed method, which integrates prototypical [15] and supervised contrastive learning [16]. The framework consists of two key components: 1) Supervised Contrastive Embedding (SCE) learns discriminative feature representations by minimizing distances between feature vectors of semantically similar regions (e.g., foreground-foreground or background-background pairs) while maximizing distances between dissimilar regions (foreground-background pairs). This ensures that semantically related image regions cluster in the embedding space. 2) Prototypical Contrastive Loss (PCL) enhances feature separation by attracting foreground samples to class-specific prototypes and repelling background samples from these prototypes. This reduces the training burden by focusing on distinguishing foreground from background, which often exhibits high intra-class variability. In Addition, the model identifies and prioritizes challenging background samples based on their distance from foreground prototypes, enhancing the discriminative capability and robustness of the learned representations.
@@ 3.3 Architecture overview and loss function
The MSA-UNet3+ architecture extends the U-Net3+ framework for medical image segmentation, specifically optimized for coronary DSA analysis. As shown in Fig. 2, the model employs an encoder-decoder structure with three key innovations: First, the encoder performs hierarchical downsampling to extract multi-scale features, while the bottleneck layer integrates an Atrous Spatial Pyramid Pooling (ASPP) module for multi-context feature aggregation. Second, Contextual Attention Fusion Modules (CAFM) combine encoder outputs through multi-feature fusion, effectively merging fine-grained details with high-level semantic context. This dual-scale integration enhances segmentation performance for structures of varying sizes and complexity. Finally, the decoder progressively reconstructs the segmentation mask through upsampling and feature recombination with skip connections. The architecture generates two outputs: (1) a segmentation mask through the final convolutional layer, and (2) a compact feature representation via an embedding layer for auxiliary analysis. Together, these design elements collectively enable MSA-UNet3+ to achieve state-of-the-art performance in challenging coronary DSA segmentation tasks, particularly for complex anatomical structures affected by noise and low contrast.
Since its introduction, U-Net has served as the foundational architecture for medical image segmentation. Subsequent advances, including U-Net++ and U-Net3+, have significantly improved this baseline. Among these, U-Net3+ has demonstrated particular effectiveness through its innovative encoder-decoder structure that features both interconnections and intraconnections. This design preserves spatial information while capturing comprehensive multi-scale features, addressing two key limitations of U-Net++: feature insufficiency and computational inefficiency. using the preceding feature maps fully, U-Net3+ achieves superior learning efficiency with reduced computational overhead. Building on these strengths, we employ U-Net3+ as the backbone for our proposed MSA-UNet3+ architecture.
Figure 3: Architecture of the M-Encoder (Multi-Scale Attention module), highlighting the Squeeze-and-Excitation (SE) block implementation.
3.3.1 M-Encoder (Multi Scale Attention Encoder)
Traditional encoders often exhibit limitations in feature extraction, spatial information preservation, and contextual understanding. To overcome these challenges, we propose an M-Encoder inspired by [45], as illustrated in Fig. 3. The architecture employs three progressive convolutional layers that refine and concatenate features, enabling deeper and more comprehensive feature extraction. A key innovation is the integration of a Squeeze-and-Excitation (SE) block, which serves as a channel-wise attention mechanism. This component dynamically recalibrates feature responses, enhancing critical features while suppressing less informative ones. By concatenating multi-scale features prior to the SE block, the M-Encoder achieves richer feature representations at each layer. The process concludes with a dimensionality-reduction convolution that maintains essential information while improving computational efficiency. The complete M-Encoder operation is formally defined by Eq. (1).

### v3

@@ Abstract
Accurate segmentation of coronary Digital Subtraction Angiography (DSA) images is essential to diagnose and treat coronary artery diseases (CAD). Despite advances in deep learning, challenges such as high intra-class variance and class imbalance limit precise vessel delineation. Most existing approaches for coronary DSA segmentation cannot address these issues. Furthermore, existing segmentation networks’ encoders do not directly generate semantic embeddings, which could enable the decoder to reconstruct segmentation masks effectively from these well-defined features. We propose a Supervised Prototypical Contrastive Loss (SPCL) that combines supervised and prototypical contrastive learning to enhance coronary DSA image segmentation. The supervised contrastive loss enforces semantic embeddings in the encoder, improving feature differentiation. The prototypical contrastive loss allows the model to focus on the foreground class while alleviating the high intra-class variance and class imbalance problems by concentrating only on the hard-to-classify background samples. We implement the proposed SPCL loss within an MSA-UNet3+: a Multi-Scale Attention-Enhanced UNet3+ architecture. The architecture integrates key components: a Multi-Scale Attention Encoder (M-encoder) and a Multi-Scale Dilated Bottleneck (MSD-Bottleneck) designed to enhance multi-scale feature extraction and a Contextual Attention Fusion Module (CAFM) designed to preserve fine-grained details while improving contextual understanding. Experiments on a private coronary DSA dataset demonstrate that MSA-UNet3+ outperforms state-of-the-art methods, achieving the highest Dice coefficient and F1-score and significantly reducing ASD and ACD. The developed framework provides clinicians with precise vessel segmentation, enabling accurate identification of coronary stenosis and supporting informed diagnostic and therapeutic decisions. The code will be released at https://github.com/rayanmerghani/MSA-UNet3plus.
@@ 1 Introduction
Figure 1: Illustration of the desired semantic embeddings characteristics of an encoder, which should place features from the same class close together while distancing features from different classes: SCE optimizes the embedding space by minimizing the distance between similar foreground samples (in blue) and maximizing the distance between dissimilar ones. PCL focuses on learning prototypes for foreground samples (in blue star), pulling them close to their respective prototypes while pushing hard negative instances (those close to the prototypes) further away.
@@
Segmentation networks in the reviewed studies, particularly encoder-decoder architectures, face critical limitations. Encoders often prioritize performance over semantic embedding, failing to cluster similar class vectors closely in the embedding space. This challenge intensifies in coronary DSA segmentation due to the background class’s structural diversity and ambiguous boundaries, causing significant intra-class variance and class imbalance. Even current approaches focusing on semantic embedding struggle with the negative class’s extreme diversity. Additionally, most of these methods neglect hard negative samples, the most optimization-informative yet least abundant because of visual dissimilarity. To address these limitations, we propose a deep learning framework that integrates prototype learning [14] and supervised contrastive loss [15]. Our hybrid approach (Fig. 1) improves encoder discriminability by: (1) clustering positive features while distancing them from negatives via supervised contrastive loss, and (2) refining background embeddings through prototypical contrastive loss that isolates hard negatives by prototype distance. This dual strategy explicitly targets class imbalance and intra-class variance by focusing on challenging background samples.
@@
Figure 2: The architecture of the proposed MSA-UNet3+ model. The model integrates a Multi-Scale Dilated Bottleneck (MSD-Bottleneck) for multi-scale feature extraction and a Contextual Attention Fusion Module (CAFM) for enhanced contextual understanding. The M-encoder employs convolutional and transposed convolutional layers, while the decoders reconstruct the segmentation mask. This architecture enables precise segmentation of coronary arteries in DSA images by capturing both fine-grained details and broader structural information.
@@
Building on this, we integrate the proposed Supervised Prototypical Contrastive Loss (SPCL) within our Multi-Scale Attention-modified UNet3+ (MSA-UNet3+) framework (Fig. 2). The architecture enhances UNet3+ through three key modifications: (1) a Multi-Scale Attention Encoder (M-encoder) replacing the traditional encoder for hierarchical feature extraction, (2) a Multi-Scale Dilated Bottleneck (MSD-Bottleneck) with Atrous Spatial Pyramid Pooling (ASPP) for multi-contextual feature fusion, and (3) a Contextual Attention Fusion Module (CAFM) that performs channel-wise feature recalibration. The decoder then synthesizes these refined multi-scale features to generate precise coronary DSA segmentation masks.
The Key contributions of this work can be summarized as follows.
  • We propose Supervised Prototypical Contrastive Loss (SPCL), a new hybrid loss function that combines the strengths of supervised contrastive learning and prototypical contrastive learning. Unlike existing approaches, SPCL explicitly enforces ideal embedding properties in the encoder through contrastive loss, while its prototypical component targets class imbalance and high intra-class variance by dynamically focusing on hard-to-classify background samples. This dual mechanism not only improves feature discriminability but also significantly enhances segmentation accuracy—addressing two critical limitations of current methods: (1) poor embedding behavior in encoders and (2) bias toward dominant classes. Experimental results demonstrate that SPCL outperforms state-of-the-art loss functions in coronary DSA segmentation, particularly in preserving fine vessel structures and suppressing background noise.
  • Our method uniquely emphasizes and handles hard-to-classify background samples, which are generally ignored in conventional coronary DSA segmentation approaches. Unlike existing methods that treat all background regions uniformly, our framework strategically focuses on these challenging cases through prototype-guided contrastive learning. This targeted approach achieves three key advancements: (1) substantially improved learning efficiency by concentrating computational resources on diagnostically ambiguous areas, (2) more robust segmentation performance, and (3) clinically relevant handling of class imbalance that better reflects real-world coronary imaging conditions. The resulting framework shows superior generalization capability compared to current clinical implementations.
  • We present MSA-UNet3+, a novel Multi-Scale Attention-enhanced U-Net3+ architecture that offers new standards for coronary DSA image segmentation through three key innovations: (1) a hierarchical Multi-Scale Attention Encoder that adaptively focuses on vessel structures across different resolutions, (2) a Multi-Scale Dilated Bottleneck with pyramid dilation rates (1,2,4,8) for comprehensive feature extraction, and (3) Contextual Attention Fusion Modules (CAFM) that dynamically recalibrate spatial and channel-wise features. Specifically optimized for coronary DSA challenges. The architecture’s unique design enables unprecedented precision in delineating complex coronary anatomies, particularly in low-contrast regions where conventional methods typically fail.
@@ 2 Related work
Medical image segmentation is critical for computer-aided diagnosis and treatment planning [16, 17]. While U-Net and its variants [18, 19] excel at capturing spatial contextual information. However, standard U-Net struggles with fine details. Variants like Res-UNet and Attention Res-UNet address this issue. Recent advances include (1) Convolutional Neural Networks (CNNs)-based architectures for robust feature extraction [17, 20, 21], (2) Fully Convolutional Networks (FCNs) for flexible input sizes [5], (3) Recurrent Neural Networks (RNNs) / ConvLSTMs for temporal modeling [22, 23], and (4) Generative Adversarial Networks (GANs) for data augmentation [24]. Transformers (e.g., TransUNet [25]) and foundation models such as SAM [19] now enable attention-based long-range dependencies. However, class imbalance remains a fundamental challenge, particularly in medical imaging where background dominance biases model performance [26].
@@ 2.1 Traditional approaches to class imbalance
Traditional methods for handling class imbalance in machine learning, such as data resampling techniques, can be applied to medical image segmentation. Oversampling, such as creating synthetic samples from the minority class, aims to balance the class distribution [27]. However, naive oversampling methods can lead to overfitting, especially in medical images, where generating truly representative synthetic data is challenging. Undersampling, on the other hand, involves removing samples from the majority class [28]. Although this can reduce computational costs and mitigate overfitting, it leads to information loss and potentially discarding valuable data. The effectiveness of these techniques is highly dependent on the specific dataset and the nature of the class imbalance [28]. However, even with oversampling, careful consideration must be given to the methods used to generate synthetic data to ensure that they accurately represent the underlying data distribution and prevent overfitting [27].
@@ 2.2 Loss function modifications
Addressing class imbalance directly with the loss function is another common strategy. While standard cross entr- opy is sensitive to class imbalance and thus favors majority classes [29, 30], three key adaptations exist: (1) Weighted cross-entropy rebalances class contributions through manual weighting, preventing the model from being dominated by the majority class [30, 31], However, determining appropriate weights can be challenging and requires potentially hyperparameter tuning [32]; (2) Focal loss [29, 33] reduces the contribution of easily classified samples (often from the majority class) and focuses learning on more challenging samples from the minority classes; and (3) Tversky loss [34], a generalization of the Dice loss, offers another avenue to address class imbalance by adjusting the weights assigned to false positives and false negatives, allowing for a more balanced trade-off between precision and recall. Selection between losses depends on the dataset characteristics and desired precision-recall tradeoffs [29].
@@ 2.3 Advanced techniques
Advanced techniques beyond traditional and loss function-based approaches address the class imbalance in medical image segmentation. Semi-supervised learning, using labeled and unlabeled data, shows promise [35, 36]. Methods like contrastive learning and consistency regularization [37, 38] learn robust representations from limited labeled data, enhancing minority class performance. These approaches involve data augmentation and enforcing consistency between predictions on augmented versions of the same image. For example, the multitask contrastive learning framework addresses class imbalance through global and local contrastive learning with multi-scale uncertainty estimation [37]. Additionally, the dual-path framework tackles intra-class imbalance by separating objects into subclasses and using weighted maps to guide segmentation [39].
Although most contrastive learning studies focus on unsupervised settings, Khosla et al. [40] extended it to supervised learning by treating all samples from the same class as positive. Contrastive learning has also been applied to semantic segmentation. Chaitanya et al. [41] introduced techniques to exploit structural similarities in volumetric medical images within the self-supervised framework, categorizing slices of similar volumes as positive pairs and dissimilar ones as negative pairs. Similarly, Zhao et al. [42] proposed a fine-tuning approach using contrastive loss for semantic segmentation. However, these studies primarily aim to improve performance with limited data, rather than training segmentation encoders to inherently learn semantic embeddings. Despite significant progress, challenges remain in addressing class imbalance in medical image segmentation. Generating high-quality synthetic data is difficult, and the optimal choice of loss functions or augmentation strategies is often dataset-specific [43].
Overcoming these challenges is critical for accurate coronary DSA image segmentation. This study integrates supervised and prototypical contrastive learning strategies, proposing a novel hybrid loss combining supervised contrastive loss with prototypical contrastive loss. This approach enhances the encoder’s ability to generate discriminative features, enabling better differentiation between diverse classes. By focusing on challenging negative samples, hybrid loss effectively addresses class imbalance and high intra-class variance.
@@ 3.1 Overview
In this study, we first propose the Supervised Prototypical Contrastive Loss (SPCL) to resolve coronary DSA segmentation’s core challenges: high intra-class variance from ambiguous backgrounds (ribs, spine, diaphragm, and lung) and class imbalance. SPCL uniquely combines supervised contrastive learning’s discriminative power with prototype-based hard mining for class-aware embeddings. Using SPCL, we then design MSA-UNet3+, a multi-scale attention network that achieves clinically significant vessel-background separation.
Figure 3: Framework overview of MSA-UNet3+ and SPCL loss. The encoder extracts multi-scale features. Feature embeddings are jointly optimized via the Supervised Prototypical Contrastive Loss (SPCL) for discriminative representation learning and segmentation loss (Dice + BCE (Binary Cross Entropy)) for precise boundary delineation.
@@ 3.2 Supervised prototypical contrastive learning
Fig. 1 illustrates our proposed method pipeline, which integrates prototypical [14] and supervised contrastive learning [15]. The framework consists of two key components: 1) Supervised Contrastive Embedding (SCE) learns discriminative feature representations by minimizing distances between feature vectors of semantically similar regions (foreground-foreground or background-background pairs) while maximizing distances between dissimilar regions (foreground-background pairs). This ensures that semantically related image regions cluster in the embedding space. 2) Prototypical Contrastive Loss (PCL) enhances feature separation by attracting foreground samples to class-specific prototypes and repelling background samples from these prototypes. This reduces the training burden by focusing on distinguishing foreground from background, which exhibits high intra-class variability. In addition, the model prioritizes challenging background samples based on their distance from foreground prototypes, enhancing the discriminative capability. To this end, we use the resized ground-truth masks for supervision as depicted in Fig. 3
@@ 3.3 Architecture overview and loss function
U-Net remains foundational for medical image segmentation. Advances like U-Net++ and U-Net3+ improved this baseline. U-Net3+ excels via an encoder-decoder structure with interconnections and intraconnections, preserving spatial information while capturing multi-scale features. It achieves superior learning efficiency with reduced computational overhead. Building on these strengths, we employ U-Net3+ as the backbone for the proposed MSA-UNet3+.
MSA-UNet3+ extends U-Net3+ for medical image segmentation, optimized for coronary DSA. As shown in Fig. 2, it uses an encoder-decoder structure with three innovations: First, the multi scale attention encoder (M-encoder) instead of normal encoder performs hierarchical downsampling to extract multi-scale features. Second, the multi-scale dilated bottleneck (MSD-Bottleneck) layer integrates the Atrous Spatial Pyramid Pooling (ASPP) module for multi-context feature aggregation. Third, Contextual Attention Fusion Modules (CAFM) combine encoder outputs through multi-feature fusion, combining fine details with semantic context. The architecture generates two outputs: (1) a segmentation mask through the final convolutional layer and (2) a compact feature representation via an embedding layer for auxiliary analysis as in Fig Fig. 3. Together, these design elements collectively enable MSA-UNet3+ to achieve state-of-the-art performance in challenging coronary DSA segmentation tasks, particularly for complex anatomical structures affected by noise and low contrast.
Figure 4: The architecture of the M-Encoder (Multi-Scale Attention module) highlights the Squeeze-and-Excitation (SE) block implementation.
3.3.1 Multi-scale attention encoder (M-Encoder)
Traditional encoders face limitations in feature extraction, spatial preservation, and contextual understanding. We propose an M-Encoder (Fig. 4) inspired by [44], using three progressive convolutional layers to refine and concatenate features for deeper and more comprehensive feature extraction. A key innovation is integrating a Squeeze-and-Excitation (SE) block, which serves as a channel-wise attention mechanism that dynamically recalibrates channels, enhancing critical features while suppressing less informative ones. By concatenating multi-scale features before the SE block, the M-Encoder achieves richer representations at each layer, followed by dimensionality-reduction convolution for computational efficiency. The complete M-Encoder operation is formalized in Eq. (1):